OpenStudy (anonymous):

Find the polynomial function with roots 1, –2, and 5

OpenStudy (anonymous):

(x-1)(x+2)(x-5) = f(x) solve it you get your answer :)

OpenStudy (kinggeorge):

If you expand the following, you get\[(x-1)(x+2)(x-5)\]\[=x^3-4x^2-7x+10\]Thus, \(x^3-4x^2-7x+10\) is a polynomial with roots 1, -2, and 5.